OpenStudy (anonymous):

divide the polynomials. 2y^3 - 2y^2 + 3y - 9 over 2y^2

OpenStudy (lalaly):

\[\large{\frac{2y^3-2y^2+3y-9}{2y^2}}=\frac{2y^3}{2y^2}-\frac{2y^2}{2y^2}+\frac{3y}{2y^2}-\frac{9}{2y^2}\]\[=\large{y-1+\frac{3}{2y}-\frac{9}{2y^2}}\]

OpenStudy (anonymous):

thank you again, ur amazing

OpenStudy (lalaly):

your welcome:D